OpenStudy (anonymous):

when I'm solving for a measurement in an angle formed by parallel lines and transversals, how do i solve for both X and Y ?! please help

OpenStudy (anonymous):



OpenStudy (anonymous):

Alternate interior angles. Set them equal and solve for one variable. Then plug that variable in to find the other.

OpenStudy (anonymous):

ohhh okay i get it but what do you mean plug the variable in to find the other?

OpenStudy (anonymous):

@AJ17

OpenStudy (anonymous):

In this case, 2x = 3y - 2x. Solving further gives you 4x = 3y, so y = 4/3x and x = 3/4y.

OpenStudy (anonymous):

I misread the question; this is all you can do for this problem.